### v1

# SkewRoute: Training-Free LLM Routing for Knowledge Graph Retrieval- Augmented Generation via Score Skewness of Retrieved Context

Abstract
Large language models excel at many tasks but often incur high inference costs during deployment. To mitigate hallucination, many systems use a knowledge graph to enhance retrieval-augmented generation (KG-RAG). However, the large amount of retrieved knowledge contexts increase these inference costs further. A promising solution to balance performance and cost is LLM routing, which directs simple queries to smaller LLMs and complex ones to larger LLMs. However, no dedicated routing methods currently exist for RAG, and existing training-based routers face challenges scaling to this domain due to the need for extensive training data. We observe that the score distributions produced by the retrieval scorer strongly correlate with query difficulty. Based on this, we propose a novel, training-free routing framework—the first tailored to KG-RAG—that effectively balances performance and cost in a plug-and-play manner. Experiments show our method reduces calls to larger LLMs by up to 50% without sacrificing response quality, demonstrating its potential for efficient and scalable LLM deployment.
SkewRoute: Training-Free LLM Routing for Knowledge Graph Retrieval- Augmented Generation via Score Skewness of Retrieved Context
Hairu Wang1,3thanks: Equal contribution., Yuan Feng1,311footnotemark: 1, Yukun Cao1,3, Xike Xie2,3thanks: Corresponding author., S Kevin Zhou2,3 1School of Computer Science, University of Science and Technology of China, China 2School of Biomedical Engineering, USTC, China 3Data Darkness Lab, MIRACLE Center, Suzhou Institute for Advanced Reasearch, USTC, China {hrwang00,yfung}@mail.ustc.edu.cn, {xkxie,skevinzhou}@ustc.edu.cn
1 Introduction
Large Language Models (LLMs) OpenAI et al. (2024); Llama Team (2024); Qwen et al. (2025); Team et al. (2024) exhibit remarkable capabilities across a wide range of tasks. However, LLMs often hallucinate due to a lack of up-to-date or domain-specific knowledge Hong et al. (2023); Wang et al. (2023). To mitigate this, LLMs are often augmented with retrieved external knowledge Asai et al. (2024); Yu et al. (2024); Gao et al. (2024) during application. Notably, many methods use externally structured Knowledge Graphs (KGs) as external bases, retrieving query-relevant context to enhance subsequent LLM generation LUO et al. (2024); He et al. (2024); Hu et al. (2024); Li et al. (2025). KGs typically represent knowledge as triples. In Retrieval-Augmented Generation (RAG), each triple serves as a knowledge context unit, scored based on its semantic and structural relevance to the given query. Subsequently, a set of high-scoring candidates is selected and concatenated to provide the LLM with supplementary knowledge, thereby enhancing generation quality. For instance, SubgraphRAG Li et al. (2025) employs a lightweight MLP to score independent triple as knowledge context and achieves state-of-the-art performance by concatenating these contexts.
Figure 1: A Training-Free Routing Framework for LLMs in KG-RAG. Scores of retrieved contexts sorted in descending exhibit distinct skewness pattern. The framework utilizes the score skewness of retrieved contexts to route requiring no training.
Given the autoregressive nature of LLMs, inference costs are directly proportional to the number of tokens Hao et al. (2025); Snell et al. (2025). While KG-RAG mitigates hallucinations and improves generation quality, it dramatically expands input tokens due to the increased retrieved context. As shown in Figure 2(a), answering a question directly with an LLM requires 62 tokens per question. However, retrieving 100 triples using SubgraphRAG increases this to 1873 tokens per question, resulting in an over 30x increase in input token consumption. This significant overhead hinders practical LLM application.
To reduce LLM inference costs in applications, recent studies have focused on LLM routing Aggarwal et al. (2024); Ding et al. (2024); Ong et al. (2025); Feng et al. (2025), routing simple questions to smaller LLMs and complex questions to larger LLMs. This is primarily because the performance benefits of increased model scale exhibit diminishing returns; for example, Qwen2.5-14B-Instruct incurs less than twice the inference cost of Qwen2.5-7b-Instruct while delivering a 7.45% performance increase. In contrast, Qwen2.5-72b-Instruct requires almost 7x the inference cost of Qwen2.5-14b-Instruct but only achieves a 2.12% improvement, as illustrated in Figure 2(b). However, while recent routing methods have been explored in direct LLM QA applications, their application and extensibility to RAG domains is limited, primarily due to two reasons:
(a) Token vs. Context
(b) Performance vs. Cost
Figure 2: Token and Performance-Cost Statistics on CWQ. (a) illustrates how input tokens varies with retrieved contexts in KG-RAG. (b) presents inference cost and performance on LLM cloud service platform of different LLM scales.
(1) Training-Based Routing is Difficult to Construct: Existing methods rely on training a classifier to route between multiple LLMs, which inherently requires substantial amounts of high-quality labeled data specifying the optimal model for a given query. Acquiring this data via expert annotation is costly and doesn’t generalize well to new datasets, typically necessitating retraining or fine-tuning.
(2) Knowledge Source Shift Drives Routing Objective Transformation: In direct LLM QA, the LLM router relies on the internal knowledge differences between models for routing. This differs fundamentally from common RAG scenarios, where knowledge dependence is shifted to externally retrieved knowledge. This reduces the importance of inherent model knowledge and increases focus on the LLMs’ reasoning ability over external knowledge. This difference in routing objectives makes existing routing methods unsuitable for RAG.
These challenges highlight the significant hurdles of effective routing LLMs in RAG scenarios, which raises a critical question: Can we develop a training-free routing method, specifically tailored for RAG, that routes by focusing target reasoning on external knowledge sources?
By analyzing the score distribution of retrieved contexts in KG-RAG, we surprisingly find significant differences for different queries, as shown in Figure 1. The score distribution for some queries exhibits high skewness, characterized by power-law behavior,which is widely observed across many fields.  [^1] For a detailed introduce of power-law, please see Section 2. Takahashi and Tanaka-Ishii (2019); Neumann and Gros (2024); Cao et al. (2024); Feng et al. (2024). A handful of contexts dominate with high scores, leaving a long tail of many lower-scoring ones. This indicates that only a few knowledge contexts are highly relevant to the query, while the majority are not. However, other queries show a distinctly different, low skewness distribution, where a substantial portion of the contexts exhibit high relevance scores. This skewness variation is strongly correlated with the difficulty of reasoning over external knowledge. For instance, simple questions may require only simple matching within a single context, while difficult questions often demand multi-hop reasoning across multiple knowledge contexts. Thus, the varying skewness of the score distribution may serve as an indicator of query difficulty.
Inspired by the score skewness of retrieved information in KG-RAG, we first investigate the correlation between skewness patterns and query difficulty. Building on this insight, we propose the first training-free routing framework to address the existing routing challenges in KG-RAG scenarios. This framework employs a lightweight and efficient workflow, as shown in Figure 1, that leverages the skewness pattern of the retrieved context scores to dynamically route queries. Simpler queries are assigned to smaller LLMs, while more difficult queries are routed to larger LLMs, effectively balancing performance and cost. Our main contributions are summarized as follows:
  • $\bullet$
    To the best of our knowledge, this work is the first to introduce a training-free routing framework tailored for KG-RAG, leveraging the skewness patterns in retrieved scores to dynamically assess query difficulty, offering a novel perspective for future efficient RAG research.
  • $\bullet$
    We propose a novel training-free framework that fully leverages the skewness pattern of retrieved scores from external knowledge bases, enabling it to balance cost and performance between different LLMs with high flexibility.
  • $\bullet$
    Extensive evaluations across multiple LLM scales demonstrate the superiority and efficiency of our proposed routing framework, highlighting its strong generalization capability and practical value for cost-effective LLM deployment.
2 Preliminaries
Power-law. Power-law Song et al. (2024); Havrilla and Liao (2024), ubiquitous in nature, reveals an inverse relationship between scores and their descending ranks: $S(n)\;=\;\frac{C}{n^{\alpha}}$. In this formulation, $n$ denotes the rank and $S(n)$ is the corresponding score. In log-log space, $logS(n)$ against $log(n)$ approximates a straight line with slope $-\alpha$. While in linear space, it appears as a few high scores followed by a long tail of lower ones. In KG-RAG, we observe that for certain queries, the scores of retrieved contexts also follow this power-law behavior.
KG-RAG. RAG integrates contexts retrieved from external knowledge bases to enhance the accuracy and credibility of LLMs. In KG-RAG, the external knowledge base is a KG, and the retrieved contexts take the form of triples $\{\tau=(h,r,t)\}$. Existing research on KG-RAG generally involves two stages: retrieval and generation. For example, the state-of-the-art KG-RAG method, SubgraphRAG trains a scorer $\mathcal{R}$ to select a subset of triples $T=\{\tau_{1},\tau_{2},\dots,\tau_{K}\}$ by their scores $W=\{s_{1},s_{2},...,s_{K}\}$ relevant to $q$ from $G$ in the retrieval phase. And then feeds $q$, an instruction $I$, and $T$ to a generator $\mathcal{G}$ (typically an LLM) to produce the final answer: $\mathcal{G}(I,T,q)\rightarrow a$, where $a\in\mathcal{A}$. Knowledge Graph Question Answering (KGQA) is often used to evaluate KG-RAG, aiming to obtain answers from KGs given query $q$, where $q\in Q$.
LLM Route. Based on scale, LLMs are broadly categorized as larger LLMs and smaller ones. Given a query $q\in Q$, a larger LLM is formally represented as a function $F_{L}(I,T,q)=a$, mapping the query $q$ and retrieved information $T$ to an answer $a$, while a smaller LLM is denoted as $F_{S}(I,T,q)=a$. Larger LLMs deliver high quality with significant cost. In contrast, smaller LLMs are less capable but substantially reduce overhead. For queries that smaller LLMs fail but larger LLMs handle correctly, using the larger LLMs boosts performance. However, for queries both LLMs answer correctly, relying on the larger LLMs incurs a substantial extra cost. Unlike previous works Aggarwal et al. (2024); Ding et al. (2024); Ong et al. (2025) that train a binary classifier to route between smaller LLMs and larger LLMs, we design training-free router $\mathcal{L}$ on the scores $W$ of retrieved contexts for given query $q$, $\mathcal{L}:\mathcal{W}\rightarrow\{F_{S},F_{L}\}$, where $\mathcal{W}$ is the space of possible $W$ and $\theta$ is a threshold.
$a=\begin{cases}F_{S}(I,T,q)&\text{if }\mathcal{L}(W,\theta)\leq 0\\ F_{L}(I,T,q)&\text{if }\mathcal{L}(W,\theta)>0\end{cases}$
3 Methodology
3.1 Skewness of Retrieved Context Scores
(a) High Skewness
(b) Low Skewness
(c) High Skewness
(d) Low Skewness
Figure 3: Score of Retrieved Contexts in CWQ. (a)(b) are plotted in linear coordinates, while (c)(d) employ a log-log scale.
The KG-RAG method typically trains a scorer that assigns higher scores to contexts relevant to the query and lower scores to irrelevant ones. Based on these scores, the top-$K$ contexts are selected to enhance LLM generation, making the quality of this selection directly affect the final answer quality. In reality, queries can be categorized as simple or difficult Berant et al. (2013); Yih et al. (2014); Min et al. (2019); Saxena et al. (2020). Simple queries often require only single-hop information matching within one knowledge context to find the answer, while difficult queries usually need multi-hop reasoning across multiple knowledge contexts. To explore this, we visualize the score distributions assigned by the scorer across different queries and find significant differences that strongly correlate with the query difficulty.
Figure 3 presents two typical examples of the top-100 context score distributions for different queries, showing clear and significant differences. Power-law distribution is highly skewed with long-tail. In Figure 3(a) and Figure 3(c), the distribution exhibits power-law behavior and shows power-law behavior: only a few contexts are relevant to the question, while most score below 0.1 and are irrelevant. By contrast, Figure 3(b) and Figure 3(d) show lower skewness, with a large portion of contexts clearly relevant—100% contexts have scores above 0.1. This difference in skewness intuitively corresponds well with the query difficulties: difficult queries naturally require reasoning across more knowledge contexts, resulting in lower skewness.
3.2 Correlation Between Score Skewness and Query Difficulty
Retrieved context scores vary in skewness across queries. To measure it, a natural metric is the area under the scores. Highly skewed distributions, where scores drop off rapidly and the low-score long tail dominates, results in smaller areas. In contrast, low skewness distributions with a wealth of high-score triples and a gentle initial descent, yield a much larger area under the scores. Concretely, we first min-max normalize the raw scores to $[0,1]$, eliminating the magnitude differences of raw scores across different queries. And then sum these normalized scores to approximate the area. Using this metric, the area computed in Figure 3(c) is 1.07, whereas in Figure 3(d) it is 65.65.
In KG-RAG, the retrieval quality directly impacts the answer generation of LLMs. To assess query difficulty, we use a straightforward metric: we check whether the top-ranked context in retrieved set contains the answer for given query. Simple query requires only a single knowledge context to answer. If the Top-$1$ contains real answer, we consider this query is simple. In this case, the scorer has accurately identified the context critical for reasoning with high confidence.
Next, we investigate the relationship between score skewness of retrieved contexts and query difficulty. To this end, we extract queries whose answers appear in the top-scoring context and compute the areas under their scores. Smaller areas correspond to higher skewness. We partition the area from their minimum to maximum into several contiguous intervals, each corresponding to a distinct level of skewness. Figure 12 illustrates the count of simple queries across different skewness levels. Overall, a consistent trend was observed in both datasets, which queries are predominantly concentrated at higher skewness levels. This allows us to select simple queries based on their skewness. The finding reveals the strong correlation between query difficulty and score skewness, which can serve as effective indicator for routing LLMs to balance performance and cost. Further validation results is provided in the Appendix A.3.3.
(a) WebQSP Dataset
(b) CWQ Dataset
Figure 4: Simple Query Number Across Skewness.
3.3 Using Skewness to Route
Data: Dataset $\mathcal{D}$, Threshold $\theta$;
1 Initialize list of answers $\mathcal{A}\leftarrow[\ ]$;
2 foreach query $q$ in Dataset $\mathcal{D}$ do
3       if $\mathcal{L}(W,\theta)$ $\leq 0$ then
4             $ans\leftarrow$ $F_{S}$($I,T,q$);
5
6      else
7             $ans\leftarrow$ $F_{L}$($I,T,q$);
8
9      Add $ans$ to $\mathcal{A}$;
10
11return $\mathcal{A}$;
Algorithm 1 Routing Framework
In this section, we present our training-free routing framework, which can be seamlessly integrated into KG-RAG in a plug-and-play manner, as shown in Algorithm 1. By leveraging the strong correlation between score skewness of retrieved contexts and query difficulty, we require no additional training. And any metric for measuring distribution skewness is compatible with our framework for LLM selection. Queries with high score skewness are routed to economical smaller LLMs, while those with low score skewness are directed to more powerful larger LLMs, thereby striking an optimal balance between response quality and inference cost. As previously mentioned, area under scores can serve as a metric of distribution skewness, yet it still has important drawbacks. First, it is highly sensitive to min-max normalization: differences in raw score ranges across queries lead to inconsistent scaling and unstable area comparisons. Second, it collapses the entire scores into a single value, sacrificing crucial information about its distribution shape. The consequence is that widely divergent distributions can produce similar area, undermining its ability to differentiate distinct skewness. To address these shortcomings, we further explore new solutions to measure score skewness. We propose a cumulative threshold-based routing method and analyze scores from the perspective of entropy and gini coefficient to make routing decisions. Compared with the straightforward area under scores, these methods offer superior and robust discriminative power between simple and difficult queries.
Cumulative Threshold-based Routing In probability and statistics, the Cumulative Distribution Function (CDF) Kolmogoroff (1933); Kolmogorov (2018) quantifies the probability that a random variable $X$ is less than or equal to a value $x$, offering a comprehensive view of the distribution. Queries with high score skewness due to the rapid decay rate, require only a few contexts to reach a cumulative probability. Conversely, queries with low score skewness necessitate a larger number of contexts to achieve the same value. Thus, cumulative threshold-based routing could guide LLM selection. The contexts need to rank in a descending order and then normalize by the score sum into a probability distribution: $p_{i}=s_{i}/\sum_{j=1}^{K}s_{j}$. Building upon this, the cumulative sum $C_{k}$ of them is computed by: $C_{k}=\sum_{i=1}^{k}p_{i}$. For a given probability $P$ (e.g., 95%), we find the smallest $k$ satisfying $C_{k}\geq P$. Queries are classified as simple if $k\leq\theta_{k}$, otherwise difficult.
Entropy-based Routing Entropy Shannon (1948) can quantify the uniformity of a probability distribution. Queries with low score skewness, characterized by with more uniform and slower decaying context scores, exhibit higher entropy, while highly skewed distributions correspond to lower entropy. Thus, entropy can effectively assess query difficulty based on context scores, enabling dynamic routing of queries. Calculating the entropy for scores requires converting them into a probability distribution $p=\{p_{1},p_{2},\dots,p_{K}\}$ by normalizing the scores $p_{i}=s_{i}/\sum_{j=1}^{K}s_{j}$. The entropy is then calculated as $H=-\sum_{i=1}^{N}p_{i}\log_{2}(p_{i})$. Thereby, a query is routed as simple if the entropy of context scores falls below a predetermined threshold $\theta_{H}$, and as difficult otherwise.
Gini Coefficient-based Routing The Gini coefficient Gates (1909); Lorenz (1905), traditionally measuring income inequality, can effectively reflect the imbalance of context scores. A higher Gini coefficient indicates greater skewness of context scores, reflecting a more pronounced imbalance among the retrieved knowledge contexts. Thus, Gini coefficient can serve as an effective metric for routing decision. The calculation of the Gini coefficient for context scores $W$ involves sorting them in ascending order $s^{\prime}_{1}\leq s^{\prime}_{2}\leq\dots\leq s^{\prime}_{K}$ and applying the following formula: $\frac{1}{K}\left(K+1-2\frac{\sum_{i=1}^{K}(K-i+1)s^{\prime}_{i}}{\sum_{j=1}^{K}s^{\prime}_{j}}\right)$. Queries are routed as simple if the calculated Gini coefficient exceeds a threshold $\theta_{G}$, indicating high skewness, and as difficult if it falls below $\theta_{G}$.
4 Experiments
4.1 Settings
(a) Qwen Hit@1
(b) Qwen F1-Score
(c) Llama Hit@1
(d) Llama F1-Score
Figure 5: Results of Routing on WebQSP. (a)(b) between Qwen7b/72b and (c)(d) between Llama8b/70b
(a) Qwen Hit@1
(b) Qwen F1-Score
(c) Llama Hit@1
(d) Llama F1-Score
Figure 6: Results of Routing on CWQ. (a)(b) between Qwen7b/72b and (c)(d) between Llama8b/70b
Datasets. To test the effectiveness of routing on knowledge-intensive reasoning tasks, we used two widely recognized KGQA datasets: CWQ Talmor and Berant (2018) and WebQSP Yih et al. (2016). Freebase Bollacker et al. (2008) serves as the background KG for both datasets, comprising more than 120 million triples. Query difficulty, assessed by the number of required knowledge contexts, varies across datasets. Specifically, WebQSP is relatively simple, containing queries with only 1 or 2 knowledge contexts. In contrast, CWQ contains more difficult reasoning problems involving up to 4 contexts. Further details of the datasets are described in Appendix A.1.
Baselines. SubgraphRAG, the latest SOTA method in KG-RAG, achieves effective and efficient retrieval of a suitable amount of relevant contexts from KGs for LLM generation. Consequently, all experiments conducted in this study are based on it. We consider three baselines: all-larger-LLM strategy, all-smaller-LLM, and random mixing.
Reasoning LLMs. We utilize a diverse selection of open-source models for our experiments: Qwen2.5-7b-Instruct, Qwen2.5-72b-Instruct Qwen et al. (2025), Llama3.1-8b-Instruct, and Llama3.1-70b-Instruct Llama Team (2024). For brevity, we will refer to them as Qwen7b, Qwen72b, Llama8b, and Llama70b, respectively. Based on their parameter counts, Qwen7b and Llama8b are classified as smaller LLMs, while their counterparts, Qwen72b and Llama70b, fall under the category of larger LLMs. Generally, larger LLMs with more parameters, deliver stronger inference performance but incur greater overhead.
Evaluation Metrics. Consistent with prior research, we employ Hit@1  [^2] Here, we use the corrected evaluation code of SubgraphRAG. In the paper, the original Hit@1 of Llama3.1-8b-Instruct with 100 triples is 56.98. and F1-Score as evaluation metrics. Hit@1 measures the rate at which the correct answer appears at the top prediction position. Given that a query may have multiple answers, F1-Score is used to assess the comprehensive coverage of all correct answers, effectively balancing precision and recall. The objective of LLM routing is to balance performance and cost, so the final evaluation metric is the performance at a specified larger LLM call ratio, equivalently Hit@1 or F1-Score under a given inference cost.
Experiment Implementations. For reasoning, we use one-shot prompting to LLMs’ generation. Consistent with SubgraphRAG, the prompt template we employe is detailed in Appendix A.4. The temperature parameter is set to 0 for reproduction, and the maximum token length for generation is fixed at 4000.
4.2 Results
Figure 5 present the experimental results of our four routing methods on the WebQSP dataset. Overall, all skewness indicators in our training-free framework consistently outperform the random mixing baseline across all scenarios on Hit@1 and F1-Score. Although the naive area-based method is less effective than other methods, its performance continuously remains above the baseline. Specifically, when routing between Qwen7b and Qwen72b, our methods cut larger LLM call ratio by half while achieving Hit@1 and F1-Score on par with full inference of larger LLM. Beyond this point, they even surpass larger LLM-only performance. Furthermore, we also test with Llama8b and Llama70b. Exemplified by the cumulative threshold-based routing method, a performance gain of nearly 2% in both Hit@1 and F1-score can be observed compared to random mixing at larger LLM call ratio nearly 0.2. These results strongly demonstrate the effectiveness of our routing methods based on score skewness to optimizing performance and cost. Additionally, we observe that the performance gap between Qwens is smaller than that between Llamas. While routing is effective in both cases, the narrower gap produces superior routing outcomes.
Figure 6 illustrates our experimental results on the CWQ, which is a more challenging KGQA dataset. Overall, our training-free methods possess and strong flexibility and generalization, performing well across various evaluation metrics and model combinations in new data scenario. Gini coefficient-based, entropy-based, and cumulative threshold-based routing methods deliver comparable performance and outperform the naive area-based method, which suffers from scaling sensitivity and capturing distribution feature. For instance, in Figure 6(c), within the larger LLM call ratio range from 0.3 to 0.6, the area-based method already exceeds the baseline, while the other methods achieve 1% higher performance than it. This underscores the importance of effectively capturing distribution characteristics.
Figure 7: Results of Routing Between Multiple Models on CWQ Dataset.
(a) WebQSP Hit@1
(b) WebQSP F1-Score
(c) CWQ Hit@1
(d) CWQ F1-Score
Figure 8: Results of Routing Between Qwen7b and Llama70b. (a)(b) on WebQSP and (c)(d) on CWQ.
4.3 Analysis
4.3.1 Routing Across Diverse Model Sizes
In this section, we further extend our methods to include an additional model size to demonstrate its generalization. Specifically, we use Qwen7b as the small LLM, Qwen72b as the large LLM, and Qwen14b as the medium LLM, which sits between them in terms of parameters, performance, and cost. The experimental results are presented in Figure 7. Notably, including the medium LLM significantly improves the trade-off between performance and cost. Routing with the medium model yields much better results than routing directly between the small and large models. Overall, our method also demonstrates clear effectiveness, consistently outperforming the random mixing baseline. This highlights the generalization and robustness of our approach, effectively enabling training-free routing across multiple LLM sizes and allowing for finer-grained cost-performance trade-offs tailored to the user’s budget.
4.3.2 Routing Across Model Families
As shown in Figure 8, we further explore routing across different model families, such as between Qwen7b and Llama70b, to evaluate the generalization ability of our method. Even when routing across distinct model families, our routing framework remains consistently effective, demonstrating strong robustness and generalization. Specifically, in terms of Hit@1 and F1-Score on the WebQSP dataset, our method achieves nearly a 3% improvement compared to the random-mixing baseline, while maintaining a cost increase of only 5%.
4.3.3 Routing Across Cumulative Probability
(a) WebQSP Hit@1
(b) WebQSP F1-Score
(c) CWQ Hit@1
(d) CWQ F1-Score
Figure 9: Results of Cumulative Threshold-based Routing with Different $P$. (a)(b) on WebQSP and (c)(d) on CWQ between Qwen7b and Qwen72b with Cumulative Probability $P$ in {0.35, 0.65, 0.95}.
We conduct an experiment to explore the impact of the probability $P$ in the cumulative threshold-based method, as shown in Figure 9. Overall, results show that our training-free methods consistently outperform the random mixing baseline for cumulative probability $P$ ranging from 0.35 to 0.95, demonstrating its robustness. For example, on the WebQSP dataset, compared to using the LLM exclusively for inference, our methods reduce cost over 50% without any loss in performance. Highly skewed distributions of contexts scores reach given $P$ with very few examples, while those with low skewness require many more. The probability choice is crucial. A low probability $P$ is easily triggered by queries with highly skewed scores, but it also allows those with low skewness to pass, thereby diminishing the router’s discriminative power. As the results illustrate, $P=0.95$ steadily outperforms $P=0.35$.
5 Related Works
In LLM routing, the necessity of maximizing quality under cost constraints is challenging. Numerous studies have sought to mitigate this issue by training a model on a specifically constructed dataset designed to determine the most suitable LLM to handle the query. For instance, AutoMix Aggarwal et al. (2024) strategically routes queries to large LLMs, employing a POMDP router based on answer confidence of small LLMs. However, it requires multiple model calls, increasing latency and overhead. Hybrid-LLM Ding et al. (2024) trains a BERT-based router that performs binary classification based on the predicted query difficulty. RouteLLM Ong et al. (2025) leverages human preference data and employs data augmentation techniques to enhance the router performance. Afterwards, GraphRouter Feng et al. (2025) builds router for LLM selections from the graph perspective and models the LLM selection problem as an edge prediction task. Compared to them, our methods offer a novel solution that does not necessitate training on specific datasets. This not only saves substantial data synthesis and training costs but also ensures high flexibility and generalization, particularly in scenarios like KG-RAG where dedicated training data is infeasible.
Knowledge Graph Question Answering (KGQA) serves as a crucial application for evaluating KG-RAG, which requires precise answers from structured KGs. These KG-RAG methods are primarily divided into two categories: decoupled retrieval methods and integrated retrieval methods based on whether LLMs dynamically participate in the retrieval process on KGs. The integrated retrieval methods Jiang et al. (2024); Sun et al. (2024); Ma et al. (2024, 2025) exemplified by ToG, enable the LLMs to interactively explore related entities and relations on KGs and perform reasoning based on the retrieved knowledge. However, this tight coupling requires frequent LLM calls, leading to substantial inference overhead. The decoupled retrieval methods Mavromatis and Karypis (2024); LUO et al. (2024); He et al. (2024); Hu et al. (2024); Wang et al. (2025); Li et al. (2025) focus on distilling prior knowledge from KGs into LLMs. For instance, RoG fine-tunes LLMs to generate relation paths as faithful plans that are then used to retrieve valid paths from KGs. However, this method incurs substantial computational expense due to fine-tuning and processing latency, which can amount to thousands of seconds per query. In contrast, SubgraphRAG, demonstrates superior effective and efficient results by integrating a lightweight multilayer perceptron (MLP) with structural and semantic information. Accordingly, all our subsequent experiments are conducted based on it.
6 Conclusion
In this paper, we present the first LLM routing framework tailored for KG-RAG in a plug-and-play manner. This is motivated by the observation that the skewness of retrieval score distributions naturally reflects query difficulty in reasoning over external knowledge, aligning well with the RAG paradigm. By exploiting this skewness, our method dynamically routes queries among LLMs of different scales, achieving an effective trade-off between performance and cost. This simple yet robust solution also overcomes prior methods’ dependence on costly training data and generalizes well across datasets and model sizes. Extensive experiments validate the effectiveness of our approach. Our method offers a new perspective for efficient RAG deployment by reducing costs through LLM routing. Future work will focus on extending this training-free routing paradigm to broader RAG settings, including text-based and multimodal scenarios.
Limitations
SkewRoute is a highly flexible and generalizable method based on score skewness. However, current experiments focus mainly on KG-RAG, while its principles should also apply to broader RAG settings, such as chunk-based text RAG or multimodal RAG. Consequently, exploring SkewRoute’s effectiveness across these wider RAG scenarios is a promising direction. We acknowledge this as a limitation of the current work and leave it for future study.
Appendix A Appendix
A.1 Datasets
The WebQuestionsSP (WebQSP) dataset, extends the original WebQuestions dataset by providing 4,737 natural language questions annotated with corresponding SPARQL queries executable against Freebase. The ComplexWebQuestions (CWQ) dataset is designed to assess models’ abilities to answer complex queries that necessitate reasoning across multiple knowledge contexts of KGs. WebQSP and CWQ are widely utilized for training and evaluating KG-RAG methods.
Datasets Train Test Max hop
CWQ 27,639 3,531 4
WebQSP 2,826 1,628 2
Table 1: Statistics of Datasets.
Table 1 summarizes the statistics of the datasets used in our experiments. For simplicity, we define one knowledge context critical for reasoning as a single hop. Specifically, the WebQSP dataset consists of 2,826 training samples and 1,628 test samples, with a maximum hop count of 2. The CWQ dataset contains 27,639 training samples and 3,531 test samples, with a maximum hop count of 4.
Dataset 1 hop 2 hop $\geq$ 3 hop
WebQSP 65.5 % 34.5% 0.00%
CWQ 40.9 % 38.3% 20.8%
Table 2: Statistics of Query Hops of Datasets
Table 2 details the distribution of query hops across the datasets. The WebQSP dataset is relatively simpler, comprising only 1-hop and 2-hop queries, with 1-hop queries forming the majority. In contrast, CWQ is a more complex dataset. Although it comprises 1-hop and 2-hop queries, a considerable percentage (20.8%) consists of queries requiring three or four hops.
A.2 Performance and Cost
The performance of the KG-RAG state-of-the-art method SubgraphRAG on the CWQ and WebQSP datasets is presented in Table 3 with 100 retrieved triples.
Models Llama-8b Llama-70b Qwen-7b Qwen-72b
CWQ
F1-Score 46.83 53.53 42.77 52.11
Hit@1 49.90 57.94 45.68 55.25
WebQSP
F1-Score 69.29 73.93 67.55 70.76
Hit@1 78.56 84.15 77.52 80.84
Table 3: Results of SubgraphRAG with 100 Triples on CWQ and WebQSP datasets.
All our experiments are conducted on the LLM cloud service platform. Table 3 presents the cost per one million tokens of different LLMs. More pricing details are provided on the official Silicon Flow website [^3] https://siliconflow.cn/.
Models Qwen7b Qwen14b Qwen32b Qwen72b Llama8b Llama70b
Cost/M tokens 0.0485 0.0970 0.1746 0.5724 0.0485 0.5724
Table 4: Inference Cost (\$) of LLMs with Different Scale on SiliconFlow.
Driven by architectural innovations and scaling laws, LLMs has emerged. As shown in the Table 3 and Table 3, the cost-efficient smaller LLMs have limited capabilities while powerful larger LLMs incur high inference cost. This inherent performance-cost trade-off poses a significant deployment challenge. LLM routing offers an effective solution by directing queries to balance response quality with cost efficiency. For users, routing enables them handling simpler queries on their devices locally while call to expensive cloud APIs only for more difficult queries. For LLM platform providers, routing facilitates automatically directing simple queries to lower-cost smaller LLMs on the backend without sacrificing user experience. Consequently, LLM routing provides a flexible and cost-effective method, helping to unlock the full potential of LLMs across various scales while accommodating diverse budget and performance requirements.
A.3 Analysis Experiments
A.3.1 Score Skewness
Figure 10 illustrates the scores of retrieved contexts for queries within the WebQSP dataset, a trend likewise observed in the CWQ dataset. Specifically, the left panel displays high skewness, with retrieved context scores declining sharply as rank increases. This highly skewed scores indicate that a few of high-ranked contexts contribute the vast majority of the total score, while the extensive long tail comprises numerous contexts with significantly lower individual scores. In stark contrast, the right panel shows low skewness, deviating from this linear drop pattern. A prominent inflection point in its upper-right region indicates that contexts, particularly those of higher scores, diminish at a substantially more gradual pace than that of high skewness.
(a) High Skewness
(b) Low Skewness
Figure 10: Score of Retrieved Contexts in WebQSP.
A.3.2 Cumulative Frequency Distribution
Building upon the previously observed variations in score skewness across queries, we visualized their cumulative frequency distributions on the WebQSP and CWQ dataset, as shown in Figure 11. These plots reveal two primary types of cumulative behavior: high skewness and low skewness. For the former, where scores are concentrated among just a few retrieved contexts, the cumulative frequency rises steeply. Thus, a high cumulative probability is achieved with only a minimal number of contexts (e.g., Figure 11(a) illustrates reaching 95% cumulative probability with just 5 triples). In contrast, the low skewness where scores are more evenly distributed among more contexts, exhibit a slower rise in cumulative frequency, requiring more to reach the same probability (e.g., Figure 11(b), 95% requires 63). Similar behavior was also observed in the CWQ dataset. This significant variation in the number of contexts required for a given cumulative probability directly reflects the differences in underlying distribution skewness of scores. This observation inspires our cumulative threshold-based method, which leverages the context count needed to achieve a preset cumulative probability to inform dynamic routing decisions.
(a) High Skewness
(b) Low Skewness
(c) High Skewness
(d) Low Skewness
Figure 11: Cumulative Frequency Distribution of Context Scores. (a)(b) Results on the WebQSP Dataset (c)(d) Results on the CWQ Dataset.
A.3.3 Correlation Analysis
To empirically validate the potential association between score skewness and query difficulty, we execute the following experimental analyses. Score skewness can be characterized using multiple metrics: area under scores, gini coefficient, entropy, and the number of contexts required to accumulate to a specified probability (e.g., 95%). High skewness typically has a smaller area, larger gini coefficient, lower entropy, and require fewer contexts for a high cumulative probability. Conversely, low skewness generally has a larger area, smaller gini, higher entropy, and require more contexts for the same threshold. Query difficulty is measured by the position of the correct answer within the descendingly sorted top-$K$ contexts. In KG-RAG, the answer position is a crucial indicator of retrieval quality. A lower answer position indicates the true answer is located at the head of the retrieved contexts, reflecting higher retrieval accuracy and facilitating LLMs’ generation. Conversely, late ranks may indicate insufficient discrimination of retriever or missing crucial context, increasing generation difficulty.
(a) Area
(b) Cumulative Threshold
(c) Entropy
(d) Gini Coefficient
Figure 12: Correlation Between Answer Position and Score Skewness. The subfigures show results for different skewness calculation metrics: (a) Area, (b) Cumulative Threshold, (c) Entropy, and (d) Gini Coefficient. Boxes span the 25th-75th percentiles with the median line. The red line shows the mean answer position of each group. ANOVA tests whether the mean answer positions differ significantly across the four groups.
The dataset is partitioned into four mutually exclusive groups based on each metric’s quartiles (Q1: 25th percentile, Q2: median, Q3: 75th percentile). The ANOVA indicates that mean answer positions differ significantly across the four groups. And we can clearly observe the trend in answer’s position in the retrieved contexts with score skewness. Specifically, as score skewness increases, the answer tends to appear at lower ranks (later positions) within the contexts. This indicates that simpler queries are associated with higher score skewness, whereas more difficult queries correlate with lower skewness, which confirms the strong correlation between query difficulty and the retrieved score skewness. Comparing area and gini coefficient (Figure 12(a) and Figure 12(d)), we find that the mean answer position varies more substantially with the gini coefficient than with area across groups, indicating that the gini coefficient offers greater sensitivity to score skewness. This suggests gini coefficient may more effectively to highlight difficulty of queries, which confirmed by our experimental section.
A.4 Prompts
The following is the detailed prompt template used in SubgraphRAG for all experiments. Due to the effects of positional encoding in the self-attention mechanism that tokens appearing later in the sequence tend to receive higher attention weights, we place triples in ascending order.
SubgraphRAG Prompt System: Based on the triplets retrieved from a knowledge graph, please answer the question. Please return formatted answers as a list, each prefixed with "ans:".   User: Triplets:
(m.011zsc4_, organization.leadership.organization, San Francisco Giants)
(m.0crtd80, sports.sports_league_participation.league, National League West)
(San Francisco Giants, time.participant.event, 2014 Major League Baseball season)
(San Francisco Giants, time.participant.event, 2012 Major League Baseball season)
(AT&T Park, location.location.events, 2010 World Series)
(San Francisco Giants, sports.professional_sports_team.owner_s, Bill Neukom)
(San Francisco Giants, time.participant.event, 2010 Major League Baseball season)
(San Francisco Giants, sports.sports_team.championships, 2010 World Series)
(San Francisco Giants, time.participant.event, 2012 World Series)
(Crazy Crab, sports.mascot.team, San Francisco Giants)
(San Francisco Giants, time.participant.event, 2010 World Series)
(San Francisco Giants, sports.sports_team.championships, 2012 World Series)
(San Francisco Giants, sports.sports_team.team_mascot, Crazy Crab)
(San Francisco Giants, sports.sports_team.championships, 2014 World Series)
(Lou Seal, sports.mascot.team, San Francisco Giants)
Question: What year did the team with mascot named Lou Seal win the World Series?   Assistant: To find the year the team with mascot named Lou Seal won the World Series, we need to find the team with mascot named Lou Seal and then find the year they won the World Series. From the triplets, we can see that Lou Seal is the mascot of the San Francisco Giants. Now, we need to find the year the San Francisco Giants won the World Series. From the triplets, we can see that San Francisco Giants won the 2010 World Series and 2012 World Series and 2014 World Series. So, the team with mascot named Lou Seal (San Francisco Giants) won the World Series in 2010, 2012, and 2014. Therefore, the formatted answers are: ans: 2014 World Series ans: 2012 World Series ans: 2010 World Series   User: Triplets: Question: